OpenStudy (anonymous):

please help (x+3)/2 = 5 (y+5)/2 = -11

OpenStudy (anonymous):

Oops.. too tired Well ehh.. \[ \frac{x+3}{2}=5 \quad //⋅2 \\ x+3=10 \quad //−3 \\ x=7 \] And.. \[ \frac{y+5}{2}=−11 \quad //⋅2 \\ y+5=−22 \quad //−5 \\ y=−27 \]

OpenStudy (cggurumanjunath):

everything is fine but what is //.2,-3,-5 all about @pitamar ? can u explain.

OpenStudy (anonymous):

Those are just the operations made on both sides of the equations, shown in the next line. Since the whole thing seems to be (unless i'm missing something) quite simple, I thought it would be better to make it clear. That's how we used to write it in here back then thought it would be more common. and eh.. it's not .2 it's x2 just with cdot thing

OpenStudy (cggurumanjunath):

ohh very good @pitamar .

OpenStudy (anonymous):

@pitamar Thanks for helping me out. it was kinda hard to understand but i think i get it now.

OpenStudy (anonymous):

Well, if it was hard to understand then it missed the point, how could I make it any clearer then?

OpenStudy (anonymous):

@pitamar when you put the (//) it was hard to make out what that was but everything else i get.

OpenStudy (cggurumanjunath):

@2cor57 Those are just the operations made on both sides of the equations, shown in the next line.